can someone help me please
How much energy, Q, is released by an earthquake with a Richter rating, R, of 8.7? Use the function R = log[Q] - 4.4

Answer by ankor@dixie-net.com(22740)

How much energy, Q, is released by an earthquake with a Richter rating, R, of 8.7?
Use the function R = log[Q] - 4.4
log(Q) - 4.4 = 8.7
log(Q) = 8.7 + 4.4
log(Q) = 13.1
Find the anti-log of 13.1
Q = 1.26(10^13) joules of energy